# Statistics of correlations in nonlinear recurrent neural networks

### v1

@@ Abstract
The statistics of correlations are central quantities characterizing the collective dynamics of recurrent neural networks. We derive exact expressions for the statistics of correlations of nonlinear recurrent networks in the limit of a large number $N$ of neurons, including systematic $1/N$ corrections. Our approach uses a path-integral representation of the network’s stochastic dynamics, which reduces the description to a few collective variables and enables efficient computation. This generalizes previous results on linear networks to include a wide family of nonlinear activation functions, which enter as interaction terms in the path integral. These interactions can resolve the instability of the linear theory and yield a strictly positive participation dimension. We present explicit results for power-law activations, revealing scaling behavior controlled by the network coupling. In addition, we introduce a class of activation functions based on Padé approximants and provide analytic predictions for their correlation statistics. Numerical simulations confirm our theoretical results with excellent agreement.
@@ 1 Introduction
Correlations of neural activity are one of the main tools we have to study the structure and function of the nervous system. They provide crucial information by measuring the statistical interdependencies between the activity of different neurons or neural regions over time [1]. The interpretation of the correlations is not easy because they are not only controlled by direct interactions between the neurons but also by the global dynamical state of the system. For instance, theoretical analysis reveal that if the network is in an asynchronous state the cross-correlations are smaller than the autocorrelations by a factor $1/N$, NN being the network size [2]. This regime has been found to be present in a wide variety of systems [3, 4, 5, 6] under quite general conditions.
@@
In this paper we present a theoretical analysis of a recurrent neural networks with non-linear transfer functions using path-integral methods. This will lead to a new and conceptually simple representation of the neural correlations in terms of a few collective variables. It will allow us to compute the moments of the covariance matrix including the corrections of order $1/N$ necessary to evaluate the participation dimension. This approach can also be extended to compute all higher order neural correlators, as well as to systematically include subleading $1/N$ corrections. Nonlinear recurrent networks have also been studied with other methods, such as dynamical mean field theory (DMFT) (see for instance [16] and references therein), and random matrix theory [17]. In fact, as we were finishing our work, the preprint [18] appeared. It uses a DMFT ansatz with random matrix techniques to study correlations of the covariance matrix of nonlinear networks. We will compare our results to theirs, finding agreement when they overlap.
@@
The work is organized as follows. In Sec. 2 we introduce the model, the path integral description and the collective fields. In Sec. 3 we evaluate the large $N$ limit via a saddle point approximation. This section contains our main result: the generating function for the connected correlations of the network. This can be used to derive all correlation functions; we give explicitly the correlators for the covariance matrix. Using this, we derive an explicit and general expression for the dimension of participation. In this section we also comment on the relation with other approaches.
@@
In Sec. 4 we present various applications of our approach. First we consider power-law activation functions, and derive analytic results for the covariance matrix statistics. Then we introduce a class of “Padé activation” functions, which have various useful theoretical properties. We compute the analytic predictions and perform detailed comparisons with numerical simulations, finding excellent agreement. We discuss $1/N$ corrections and non-equilibrium effects. Finally, Sec. 5 is devoted to the conclusions and future directions. While the main text focuses on the correlations of neural outputs, for completeness in Appendix A we show how to derive the correlations for the inputs as well.
@@ 2 Model and path integral representation
A Recurrent Neural Network (RNN) for $N$ neurons with signal variable $\phi_{i}$, $i=1,\ldots,N$, and nonlinear activation function $f(\phi)$, is described by the equation
$\tau\frac{d\phi_{i}(t)}{dt}=-\phi_{i}(t)+W_{ij}f(\phi(t))_{j}+\xi_{i}(t)\,$, (2.1)
where $\phi$ describes the input neural current, and $f(\phi)$, the activation function, gives the neural outputs. We take the convention that all repeated indices are summed up. The number of neurons $i=1,\ldots,N$ is taken to be large. We assume that the nonlinear function $f$ is regular, odd under $\phi\to-\phi$ and vanishes at the origin,
$f(\phi)=\phi+f_{1}\phi^{3}+\ldots$ (2.2)
@@
The internal noise is a random Gaussian variable with vanishing one-point function and
$\langle\xi_{i}(t)\xi_{j}(t^{\prime})\rangle=\delta_{ij}D\delta(t-t^{\prime})\,.$ (2.3)
@@
The factor of $N$ in this variance is chosen to have a well-defined large $N$ limit.
Our goal is to compute the correlation functions of the network outputs,
$\Gamma^{N_{r}}_{i_{1}\ldots,j_{1},\ldots,s_{1}\ldots}=\langle\langle f_{i_{1}}(\phi)f_{i_{2}}(\phi)\ldots\rangle_{\xi}\langle f_{j_{1}}(\phi)f_{j_{2}}(\phi)\ldots\rangle_{\xi}\ldots\langle f_{s_{1}}(\phi)f_{s_{2}}(\phi)\ldots\rangle_{\xi}\rangle_{W}\,$, (2.5)
@@
where we kept the insertion times implicit. Here $\langle\ldots\rangle_{\xi}$ denotes the average over $\xi$, $\langle\ldots\rangle_{W}$ is the average over $\xi$, and $N_{r}$ is the number of $\xi$-averages inside the $\langle\ldots\rangle_{W}$. The calculation of the correlations of the network inputs $\phi_{i}$ can be done along similar lines; this is presented in the Appendix A.
@@
It is important to notice that experimental evaluations of the correlations do not involve an average over the coupling variables but empirical averages over some sets of neurons. Here we are assuming that in the limit of large $N$ the macroscopic state does not depend on the precise value of each of their random parameters but rather only on their statistics [19]. In other words we assume the system is self-averaging. This assumption allows us to use the average over $W$ to predict the empirical averages.
@@
Consider first a single $\xi$-average. By parity, correlation functions with an odd number of insertions vanish. Correlation functions with an even number of $f$-insertions can be obtained from the partition function with a source $J_{ij}(t_{1},t_{2})$ for $f_{i}(\phi(t_{1}))f_{j}(\phi(t_{2}))$. The time dependence of the source allows to compute correlations at different times; however, while for now we will write a partition function for the general case, the main focus of this work will be on time independent (zero frequency) correlations. Introducing a Lagrange multiplier field $\tilde{\phi}_{i}$ that imposes the stochastic equation (2.1), the path-integral representation for the partition function including the $\xi$-noise average is
$Z_{\xi}(J)$ $=$ $F(W)\int D\xi\,D\phi\,D\tilde{\phi}\,e^{-\frac{1}{2D}\int_{-T}^{T}\,dt\,\xi_{i}(t)^{2}}e^{i\int_{-T}^{T}\,dt\,\tilde{\phi}_{i}(t)\left(\tau\frac{d\phi_{i}(t)}{dt}+\phi_{i}(t)-W_{ij}f(\phi(t))_{j}-\xi_{i}(t)\right)}$ (2.6)
$\qquad\times e^{\int dt_{1}dt_{2}\,J_{ij}(t_{1},t_{2})f(\phi_{i}(t_{1}))f(\phi_{j}(t_{2}))}\,$,
@@
where $F(W)$ is a normalization factor such that $Z_{\xi}(J=0)=1$. In the linear model, a short calculation gives $F(W)$ as a determinant of the $\phi$ and $\tilde{\phi}$ Green’s function, but more generally it is a complicated function that depends on the activation function and cannot be computed in closed form. We will see shortly that its effects are subleading for a large network.
@@
Now let’s turn to the generating function of $W$-averages of the form (2.5). Each $\xi$-average is represented by a generating function (2.6), so we need to introduce $N_{r}$ replicas; the replica index will be denoted by $a,b,\ldots=1,\ldots,N_{r}$, while letters $i,j\ldots$ are the physical neuron indices. Then the fields in (2.6) acquire two indices: $\xi_{i}^{a},\,\phi_{i}^{a},\,\tilde{\phi}_{i}^{a},\,f_{i}^{a}$, while $W_{ij}$ has no replica indices because we are considering a single $W$-average. Furthermore, note that by parity only correlation functions of $f_{i}^{a}$ with an even number of $i$-indices are nonvanishing. In order to generate these, it is sufficient to use a bi-source $J_{i}^{ab}$ that is diagonal in physical space and nondiagonal in replica space. For instance,
@@
$Z_{N_{r}}(J)$ $=$ $\int DWD\xi\,D\phi\,D\tilde{\phi}\,e^{-\frac{N}{2\lambda^{2}}W_{ij}W_{ij}}\,F(W)^{N_{r}}\,e^{-\frac{1}{2D}\int_{-T}^{T}\,dt\,\xi^{a}_{i}(t)^{2}}$
$\qquad e^{i\int_{-T}^{T}\,dt\,\tilde{\phi}_{i}^{a}(t)\left(\tau\frac{d\phi^{a}_{i}(t)}{dt}+\phi_{i}(t)-W_{ij}f_{j}^{a}(t)-\xi_{i}^{a}(t)\right)}e^{\int dt_{1}dt_{2}\,J_{i}^{ab}(t_{1},t_{2})f_{i}^{a}(t_{1})f_{i}^{b}(t_{2})}\,.$
The external source $J_{i}^{ab}$ is chosen to be an upper-triangular matrix in replica space. [^1] We could also choose it to be symmetric in the indices $(a,b)$, at the cost of introducing factors of $2$ between the source derivatives of $Z(J)$ and the correlation functions. As long as $N_{r}\ll N$, the factor of $F(W)^{N_{r}}$ gives a small correction to the Gaussian dynamics of $W$, and so we will neglect it in what follows.
The partition function can be written as a path integral of $e^{-S}$, with action
$S$ $=$ $\int_{-T}^{T}\,dt\left[\frac{N}{2\lambda^{2}}W_{ij}^{2}\delta(t)+\frac{1}{2D}\xi_{i}^{a}(t)^{2}-i\tilde{\phi}_{i}^{a}(t)\left(\tau\frac{d\phi_{i}^{a}(t)}{dt}+\phi_{i}^{a}(t)-W_{ij}f_{j}^{a}(t)-\xi_{i}^{a}(t)\right)\right]$ (2.9)
$-\int_{-T}^{T}dt_{1}dt_{2}\,J_{i}^{ab}(t_{1},t_{2})f_{i}^{a}(t_{1})f_{i}^{b}(t_{2})\,.$
In this paper we will focus for simplicity on equilibrium or zero-frequency correlation functions. It would be very interesting to extend the analysis below to include the time-dependent dynamics of the network, something that we plan to address in future work. In this case we can drop all the time dependence and the time integrals, the path integrals become ordinary integrals, and the action is
$S=\frac{N}{2\lambda^{2}}W_{ij}^{2}+\frac{1}{2D}(\xi_{i}^{a})^{2}-i\tilde{\phi}_{i}^{a}\left(\phi_{i}^{a}-W_{ij}f_{j}^{a}-\xi_{i}^{a}\right)-J_{i}^{ab}f_{i}^{a}f_{i}^{b}\,.$ (2.10)
@@
$Z_{N_{r}}(J)=\int d\phi^{a}_{i}\,d\tilde{\phi}^{b}_{j}\,e^{-\frac{1}{2}M_{ab}\tilde{\phi}_{i}^{a}\tilde{\phi}_{i}^{b}+i\phi_{i}^{a}\tilde{\phi}_{i}^{a}}\,e^{J^{ab}_{i}f^{a}_{i}f^{b}_{i}}$ (2.11)
@@
$M_{ab}=\delta_{ab}D+\frac{\lambda^{2}}{N}\sum_{j}\,f_{j}^{a}f_{j}^{b}\,.$ (2.12)
@@
$Z_{N_{r}}(J)=\int d\phi^{a}_{i}\;[\textrm{det}(2\pi M^{-1})]^{N/2}\,e^{-\frac{1}{2}(M^{-1})_{ab}\phi_{i}^{a}\phi_{i}^{b}}\,e^{J^{ab}_{i}f^{a}_{i}f^{b}_{i}}\,.$ (2.13)

### v2

@@ Abstract
The statistics of correlations are central quantities characterizing the collective dynamics of recurrent neural networks. We derive exact expressions for the statistics of correlations of nonlinear recurrent networks in the limit of a large number $N$ of neurons, including systematic $1/N$ corrections, in the regime of Gaussian quenched disorder. Our approach uses a path-integral representation of the network’s stochastic dynamics, which reduces the description to a few collective variables and enables efficient computation. This generalizes previous results on linear networks to include a wide family of nonlinear activation functions, which enter as interaction terms in the path integral. These interactions can resolve the instability of the linear theory and yield a strictly positive participation dimension. We present explicit results for power-law activations, revealing scaling behavior controlled by the network coupling. In addition, we introduce a class of activation functions based on Padé approximants and provide analytic predictions for their correlation statistics. Numerical simulations confirm our theoretical results with excellent agreement. We also compare with previous works that have studied the complementary case with annealed disorder, and based on this we propose a new self-consistent equation for the more general case of colored noise.
@@ 1 Introduction
Correlations of neural activity are one of the main tools we have to study the structure and function of the nervous system. They provide crucial information by measuring the statistical interdependencies between the activity of different neurons or neural regions over time [8]. The interpretation of the correlations is not easy because they are not only controlled by direct interactions between the neurons but also by the global dynamical state of the system. For instance, theoretical analysis reveals that if the network is in an asynchronous state the cross-correlations are smaller than the autocorrelations by a factor $1/N$, NN being the network size [16]. This regime has been found to be present in a wide variety of systems [1, 3, 6, 18] under quite general conditions.
@@
In this paper we present a theoretical analysis of recurrent neural networks with non-linear transfer functions and Gaussian noise, using path-integral methods. This will lead to a new and conceptually simple representation of the neural correlations in terms of a few collective variables. It will allow us to compute the moments of the covariance matrix including the corrections of order $1/N$ necessary to evaluate the participation dimension. This approach can also be extended to compute all higher order neural correlators, as well as to systematically include subleading $1/N$ corrections. Nonlinear recurrent networks have also been studied with other methods, such as dynamical mean field theory (DMFT) (see for instance [7] and references therein), and random matrix theory [36]. In fact, as we were finishing our work, the preprint [32] appeared. It uses a DMFT ansatz with random matrix techniques to study correlations of the covariance matrix of nonlinear networks.
This theoretical analysis requires modeling of the dynamics incorporating noisy contributions and the appropriate time scales. In the continuous-time recurrent neural network equation that we will use, there are two relevant time scales: the neuron time constant $\tau$, and the noise correlation time $\tau_{\textrm{noise}}$. The internal noise is often taken to be Gaussian and white, corresponding to $\tau_{\textrm{noise}}/\tau\to 0$. This is a useful mathematical simplification, but it is not always very realistic since it corresponds to an internal random variable that is changing infinitely fast. In this work we will instead take a complementary idealized limit: $\tau_{\textrm{noise}}\gg\tau$. In statistical mechanics terms, this corresponds to quenched disorder, as opposed to the annealed disorder that arises for $\tau_{\textrm{noise}}\to 0$. We will find that this provides important advantages for our goal of obtaining the large $N$ neural correlators. More realistic systems have $\tau\sim\tau_{\textrm{noise}}$, and this intermediate regime is harder to study theoretically. Nevertheless, it is interesting that the results that we obtain in the quenched limit $\tau_{\textrm{noise}}\gg\tau$ are not that different from those for $\tau_{\textrm{noise}}\ll\tau$. In this way, the quenched analysis of our work complements previous works on the white noise limit (which we will review below), and strongly suggests that the more realistic intermediate regime may after all not be that distinct from either of the mathematically simplified limits.
@@
The work is organized as follows. In Sec. 2 we introduce the model, the path integral description and the collective fields. In Sec. 3 we evaluate the large $N$ limit via a saddle point approximation. This section contains our main theoretical results: the generating function for the connected correlations of the network. This can be used to derive all correlation functions; we give explicitly the correlators for the covariance matrix. Using this, we derive an explicit and general expression for the dimension of participation.
@@
Sec. 4 contains our main numerical results, exhibiting various concrete applications of our approach. First we consider power-law activation functions, and derive analytic results for the covariance matrix statistics. Then we introduce a class of “Padé activation” functions, which have various useful theoretical properties. We compute the analytic predictions and perform detailed comparisons with numerical simulations, finding excellent agreement. We discuss $1/N$ corrections and non-equilibrium effects. We compare our results with previous ones in the literature, including work on linear networks, as well as more recent results on nonlinear RNNs in the complementary white noise regime.
Finally, Sec. 5 is devoted to the conclusions and future directions. While the main text focuses on the correlations of neural outputs, for completeness in Appendix A we show how to derive the correlations for the inputs as well. Appendix B is devoted to the numerical methods. Appendix C presents in more detail some steps of our analytical results.
@@ 2 Model and path integral representation
A Recurrent Neural Network (RNN) for $N$ neurons with signal variable $\phi_{i}$, $i=1,\ldots,N$, and nonlinear activation function $f(\phi)$, is described by the equation  [^1] Equation (2.1) is a standard continuous-time firing-rate (leaky-integrator) network driven by additive Gaussian noise, i.e. a Langevin equation. In the special linear case $f(x)=x$ (or after linearization around a fixed point), it reduces to a multivariate Ornstein–Uhlenbeck (linear Langevin) process. See e.g. [16, 12, 11, 33] for closely related rate-network dynamics and correlation theories.
$\tau\frac{d\phi_{i}(t)}{dt}=-\phi_{i}(t)+\sum_{j=1}^{N}W_{ij}f(\phi(t))_{j}+\xi_{i}(t)\,$, (2.1)
where $\phi$ describes the input neural current, and $f(\phi)$, the activation function, gives the neural outputs. The number of neurons $i=1,\ldots,N$ is taken to be large. We assume that the nonlinear function $f$ is regular, odd under $\phi\to-\phi$ and vanishes at the origin,
$f(\phi)=\phi+f_{(1)}\phi^{3}+\ldots$ (2.2)
@@
The factor of $N$ in this variance is chosen to have a well-defined large $N$ limit [^2] Let us note that this distribution is unrealistic from a biological point of view. Real weight distributions are log-normal instead of Gaussian and they also display significant correlations [34]. On the other hand, the random internal noise $\xi(t)$ can be modeled by its own Ornstein-Uhlenbeck with time-scale $\tau_{\textrm{noise}}$,
$\frac{d\xi_{i}(t)}{dt}=-\frac{\xi_{i}(t)}{\tau_{\textrm{noise}}}+\sigma\eta_{i}(t)$ (2.4)
with $\eta_{i}(t)$ a random Gaussian white noise variable, $\langle\eta(t)\eta(t^{\prime})\rangle=\delta(t-t^{\prime})$. At late times then,
$\langle\xi_{i}(t)\xi_{j}(0)\rangle\approx\delta_{ij}\,\frac{\sigma^{2}\tau_{\textrm{noise}}}{2}e^{-|t|/\tau_{\textrm{noise}}}\,.$ (2.5)
Experimentally, the timescale associated with input synchrony, which motivates our $\tau_{\textrm{noise}}$, can range from a few tens to a few hundred milliseconds, approximately 2020200200 ms, depending on whether the cortical network is in a spontaneous state or in a stimulus-driven state [28]. By contrast, intrinsic cortical neuronal timescales, related with the dynamical timescale $\tau$, are typically of order 10102020 ms [25]. The annealed (white-noise) limit corresponds to $\tau_{\textrm{noise}}\to 0$, so that the time dependence in (2.5) becomes proportional to $\delta(t)$, keeping $\sigma^{2}\tau_{\textrm{noise}}$ fixed. This limit has been analyzed before in our context by e.g. [19, 10, 32]. In this work we will instead focus on the limit $\tau\ll\tau_{\textrm{noise}}$, and dynamics with times $t$ such that $\tau\ll t\ll\tau_{\textrm{noise}}$. This means that the equilibrium configurations and their corresponding correlation functions can be obtained while keeping the random variable $\xi$ approximately constant (quenched disorder). This corresponds to taking
$\langle\xi_{i}\xi_{j}\rangle=\delta_{ij}D\,$, (2.6)
with $D\sim\sigma^{2}\tau_{\textrm{noise}}$ fixed.
2.1 Equilibrium partition function representation
We will focus for simplicity on equilibrium correlation functions. In the regime $\tau\ll\tau_{\textrm{noise}}$, the equilibrium configuration is described by
$0=-\phi_{i}+\sum_{j}\,W_{ij}f(\phi_{j})+\xi_{i}\,$, (2.7)
with $W$ and $\xi$ random quenched disorder variables as discussed before. This will provide us with important mathematical simplifications that will lead to explicit results in the large N limit. Below we will also compare with the recent results of [32] in the opposite limit $\tau_{\textrm{noise}}\ll\tau$. It would be very interesting to extend the analysis below to include the time-dependent dynamics of the network, something that we plan to address in future work. The numerical simulation of Eq. (2.1) is described in App. B.
Our goal is to compute the equilibrium correlation functions of the network outputs,
$\Gamma^{N_{r}}_{i_{1}\ldots,j_{1},\ldots,s_{1}\ldots}=\langle\langle f(\phi_{i_{1}})f(\phi_{i_{2}})\ldots\rangle_{\xi}\langle f(\phi_{j_{1}})f(\phi_{j_{2}})\ldots\rangle_{\xi}\ldots\langle f(\phi_{s_{1}})f(\phi_{s_{2}})\ldots\rangle_{\xi}\rangle_{W}\,$, (2.8)
@@ 2.1 Equilibrium partition function representation
Here $\langle\ldots\rangle_{\xi}$ denotes the average over $\xi$, $\langle\ldots\rangle_{W}$ is the average over $W$, and $N_{r}$ is the number of $\xi$-averages inside the $\langle\ldots\rangle_{W}$. The calculation of the correlations of the network inputs $\phi_{i}$ can be done along similar lines; this is presented in the Appendix A.
@@
It is important to notice that experimental evaluations of the correlations do not involve an average over the coupling variables, but rather empirical averages over some sets of neurons. Here we are assuming that in the limit of large $N$ the macroscopic state does not depend on the precise value of each of their random parameters but rather only on their statistics; see also [11] for more discussions on this point. In other words we assume the system is self-averaging. This assumption allows us to use the average over $W$ to predict the empirical averages.
@@
For the purpose of obtaining the large $N$ correlators, we will find it convenient to proceed in terms of a partition function representation. Consider first a single $\xi$-average. By the assumed parity property $f(-\phi)=-f(\phi)$, correlation functions with an odd number of insertions vanish. Correlation functions with an even number of $f$-insertions can be obtained from the partition function with a source $J_{ij}$ for $f_{i}f_{j}$. Introducing a Lagrange multiplier field $\tilde{\phi}_{i}$ that imposes the equilibrium stochastic equation (2.7), the integral representation for the generating function including the $\xi$-noise average is
$Z_{\xi}(J)$ $=\int\mathcal{D}\xi\,\mathcal{D}\phi\,\mathcal{D}\tilde{\phi}\;F(W,\xi)\;e^{-\frac{1}{2D}\sum_{i}\xi_{i}^{2}}e^{i\sum_{i}\tilde{\phi}_{i}\left(\phi_{i}-\sum_{j}W_{ij}\,f(\phi_{j})-\xi_{i}\right)}e^{\sum_{ij}J_{ij}\,f(\phi_{i})\,f(\phi_{j})}\,.$ (2.9)
@@
We denote ${\mathcal{D}}\phi=\prod_{i}d\phi_{i}$, and similarly for other integrated variables. Here $F(W,\xi)$ arises from the Jacobian of the delta function produced by the Lagrange multiplier. [^3] In other words, for the integral over all possible $\phi$, $F$ satisfies $\delta(\phi-\phi_{\textrm{sol}})=F(W,\xi)\delta(\phi_{i}-\sum_{j}W_{ij}\,f(\phi_{j})-\xi_{i})$, where $\phi_{\textrm{sol}}$ is the solution to 2.7. In the linear model, a short calculation gives $F(W,\xi)$ as a determinant of the $\phi$ and $\tilde{\phi}$ Green’s function, but more generally it is a complicated function that depends on the activation function and cannot be computed in closed form. We will see shortly that its effects are subleading for a large network.
@@
Now let’s turn to the generating function of $W$-averages of the form (2.8). Each $\xi$-average is represented by a generating function (2.9), so we need to introduce $N_{r}$ replicas; the replica indices will be denoted by $a,b,\ldots=1,\ldots,N_{r}$. Then the fields in (2.9) acquire two indices: $\xi_{i}^{a},\,\phi_{i}^{a},\,\tilde{\phi}_{i}^{a},\,f_{i}^{a}\equiv f(\phi_{i}^{a})$, while $W_{ij}$ has no replica indices because we are considering a single $W$-average. Furthermore, note that by parity only correlation functions of $f_{i}^{a}$ with an even number of $i$-indices are nonvanishing. In order to generate these, it is sufficient to use a bi-source $J_{i}^{ab}$ that is diagonal in physical space (neuron indices $i$) and nondiagonal in replica space. For instance,
@@
$Z_{N_{r}}(J)=\int\mathcal{D}W\,\mathcal{D}\xi\,\mathcal{D}\phi\,\mathcal{D}\tilde{\phi}\;F(W,\xi)^{N_{r}}\;e^{-S}\,$, (2.11)
with action
$S$ $=$ $\frac{N}{2\lambda^{2}}\sum_{i,j=1}^{N}W_{ij}^{2}+\sum_{i=1}^{N}\sum_{a=1}^{N_{r}}\left[\frac{1}{2D}\big(\xi_{i}^{a}\big)^{2}-i\tilde{\phi}_{i}^{a}\Big(\phi_{i}^{a}-\sum_{j=1}^{N}W_{ij}f_{j}^{a}-\xi_{i}^{a}\Big)\right]$ (2.12)
$\qquad-\sum_{i=1}^{N}\sum_{a,b=1}^{N_{r}}J_{i}^{ab}\,f_{i}^{a}\,f_{i}^{b}\,.$
As before, the notation $\mathcal{D}X$ stands for the products of all the differentials in the $X$ vector or matrix; so, for example $\mathcal{D}W=\prod_{ij}dW_{ij}$, $\mathcal{D}\phi=\prod_{i,a}d\phi_{i}^{a}$ and so on. The external source $J_{i}^{ab}$ is chosen to be an upper-triangular matrix in replica space. [^4] We could also choose it to be symmetric in the indices $(a,b)$, at the cost of introducing factors of $2$ between the source derivatives of $Z(J)$ and the correlation functions. As long as $N_{r}\ll N$, the factor of $F(W,\xi)^{N_{r}}$ gives a small correction to the Gaussian dynamics of $W$ and $\xi$, and so we will neglect it in what follows. [^5] More explicitly, the Jacobian factor is proportional to $|det(1-Wf^{\prime})|=|e^{\sum_{i}\log(1-W_{ii}f_{i}^{\prime})}|$. It only includes the $N$ diagonal terms. This is explicitly subleading compared to the order $N^{2}$ terms in the action for $W$, $\sum_{ij}W_{ij}^{2}$, which moreover in the large $N$ normalization has an explicit $N$ in front.
@@
$Z_{N_{r}}(J)=\int\mathcal{D}\phi\,\mathcal{D}\tilde{\phi}\;\exp\!\left[\sum_{i=1}^{N}\left(-\frac{1}{2}\sum_{a,b=1}^{N_{r}}M_{ab}\,\tilde{\phi}_{i}^{a}\tilde{\phi}_{i}^{b}+i\sum_{a=1}^{N_{r}}\phi_{i}^{a}\tilde{\phi}_{i}^{a}+\sum_{a,b=1}^{N_{r}}J_{i}^{ab}\,f_{j}^{a}\,f_{j}^{b}\right)\right]$ (2.13)
@@
$M_{ab}=\delta_{ab}\,D+\frac{\lambda^{2}}{N}\sum_{j=1}^{N}f_{j}^{a}\,f_{j}^{b}.$ (2.14)
@@
$Z_{N_{r}}(J)$ $=\int\mathcal{D}\phi\;\big[\det(2\pi M^{-1})\big]^{N/2}\exp\!\left[\sum_{i=1}^{N}\sum_{a,b=1}^{N_{r}}\left(-\frac{1}{2}(M^{-1})_{ab}\,\phi_{i}^{a}\phi_{i}^{b}+J_{i}^{ab}\,f_{i}^{a}f_{i}^{b}\right)\right].$ (2.15)